OpenStudy (anonymous):

AP Calculus: A cylindrical can is to hold a volume of 800cc. The can has double thick top. Find the dimensions to use the minimum surface area.

OpenStudy (anonymous):

so far I have S=3pir^2+2pirh 800=pir^2h, h=800/pir^2 S=3pir^2+2pir(800/pir^2) S=3pir^2+1600/r S'=6pir+-1600/r^2=0

OpenStudy (anonymous):